# Object-Centric Representation Learning for Enhanced 3D Semantic Scene Graph Prediction

### v2

@@ D.2 Quantitative Results
SGPN [wald2020learning] 96.66 99.17 66.19 85.73 10.18 28.41 15.78 29.60 66.60 77.03
SGFN [wu2021scenegraphfusion] 95.08 99.38 70.02 87.81 38.67 58.21 22.59 35.68 71.44 80.11
VL-SAT [wang2023vl] 96.31 99.21 80.03 93.64 52.38 66.13 31.28 47.26 75.09 82.25
@@
KERN [chen2019knowledge] 9.5 11.5 11.9 18.8 25.6 26.5
SGPN [wald2020learning] 19.7 22.6 23.1 32.1 38.4 38.9
SGFN [wu2021scenegraphfusion] 20.5 23.1 23.1 46.1 54.8 55.1
Zhang et al.  [zhang2021knowledge] 24.4 28.6 28.8 56.6 63.5 63.8
VL-SAT [wang2023vl] 31.0 32.6 32.7 57.8 64.2 64.3
@@
SGPN [wald2020learning] 10.86 27.54 32.01 55.22 41.52 51.92
SGFN [wu2021scenegraphfusion] 21.24 46.68 41.89 70.82 58.37 67.61
VL-SAT [wang2023vl] 21.41 46.14 54.03 77.67 65.09 73.59
@@
Performance on additional metrics. Table  9 presents the quantitative evaluation results, reporting mR@$K$ for Head, Body, and Tail predicates ($K=3,5$), along with R@$K$ for seen and unseen triplets ($K=50,100$). The proposed method demonstrates consistent performance improvements across the majority of metrics. Notably, mR@$K$ for Tail predicates—typically the most challenging category due to data sparsity—shows an improvement of more than 4% over the previous highest performing method. Additionally, R@$K$ on unseen triplets increases by approximately one percentage point. These improvements, achieved despite the pronounced long‐tailed class distribution characteristic of the 3DSSG dataset [wald2020learning], indicate enhanced generalization capabilities of the proposed approach compared to existing methods.
@@
Table 14: Extensibility to weakly-supervised settings. “w/ weakly-sup.” denotes the weakly supervised variant of our method and “w/ fully-sup.” denotes our original setting(fully-supervised manner). Under weak supervision, our approach outperforms 3D-VLAP [wang2024weakly] on both R@K and mR@K. However, it does not match our fully supervised model, likely due to the inherent noise in weakly supervised signals arising from imprecise triplet labels.
Extensibility to weakly-supervised setting. As mentioned in Section. A, weakly-supervised approach also seems feasible for 3D scene graph generation task. Following 3D-VLAP [wang2024weakly], we adopted weakly-supervised setting to our approach and check whether our hypothesis also works with other training scheme. In Table. 14, proposed method also performs well with weakly-supervised setting, overwhelming original 3D-VLAP with huge margin. This result provides us useful insights that our hypothesis also valid in other training schemes. Especially, the mR@K showed huge gap compared to 3D-VLAP, which shows promising aspects to address long-tailed distribution problem with reducing extensive human annotation labor. However, It could not exceeded fully-supervised settings suggested in manuscript. We assume that weakly-supervised siginals(pseudo-labels selected by CLIP) made errors, causing somewhat tricky to learn proper insights.
@@ D.3 Qualitative Results
Figure 7: Qualitative analysis of object feature space. (a) Class‑level cosine similarity of Ours, (b) is for same as that of VL-SAT [wang2023vl].

### v3

@@ D.2 Quantitative Results
SGPN [45] 96.66 99.17 66.19 85.73 10.18 28.41 15.78 29.60 66.60 77.03
SGFN [52] 95.08 99.38 70.02 87.81 38.67 58.21 22.59 35.68 71.44 80.11
VL-SAT [48] 96.31 99.21 80.03 93.64 52.38 66.13 31.28 47.26 75.09 82.25
@@
KERN [5] 9.5 11.5 11.9 18.8 25.6 26.5
SGPN [45] 19.7 22.6 23.1 32.1 38.4 38.9
SGFN [52] 20.5 23.1 23.1 46.1 54.8 55.1
Zhang et al.  [63] 24.4 28.6 28.8 56.6 63.5 63.8
VL-SAT [48] 31.0 32.6 32.7 57.8 64.2 64.3
@@
SGPN [45] 10.86 27.54 32.01 55.22 41.52 51.92
SGFN [52] 21.24 46.68 41.89 70.82 58.37 67.61
VL-SAT [48] 21.41 46.14 54.03 77.67 65.09 73.59
@@
Performance on additional metrics. Table  9 presents the quantitative evaluation results, reporting mR@$K$ for Head, Body, and Tail predicates ($K=3,5$), along with R@$K$ for seen and unseen triplets ($K=50,100$). The proposed method demonstrates consistent performance improvements across the majority of metrics. Notably, mR@$K$ for Tail predicates—typically the most challenging category due to data sparsity—shows an improvement of more than 4% over the previous highest performing method. Additionally, R@$K$ on unseen triplets increases by approximately one percentage point. These improvements, achieved despite the pronounced long‐tailed class distribution characteristic of the 3DSSG dataset [45], indicate enhanced generalization capabilities of the proposed approach compared to existing methods.
@@
Table 14: Extensibility to weakly-supervised settings. “w/ weakly-sup.” denotes the weakly supervised variant of our method and “w/ fully-sup.” denotes our original setting(fully-supervised manner). Under weak supervision, our approach outperforms 3D-VLAP [47] on both R@K and mR@K. However, it does not match our fully supervised model, likely due to the inherent noise in weakly supervised signals arising from imprecise triplet labels.
Extensibility to weakly-supervised setting. As mentioned in Section. A, weakly-supervised approach also seems feasible for 3D scene graph generation task. Following 3D-VLAP [47], we adopted weakly-supervised setting to our approach and check whether our hypothesis also works with other training scheme. In Table. 14, proposed method also performs well with weakly-supervised setting, overwhelming original 3D-VLAP with huge margin. This result provides us useful insights that our hypothesis also valid in other training schemes. Especially, the mR@K showed huge gap compared to 3D-VLAP, which shows promising aspects to address long-tailed distribution problem with reducing extensive human annotation labor. However, It could not exceeded fully-supervised settings suggested in manuscript. We assume that weakly-supervised siginals(pseudo-labels selected by CLIP) made errors, causing somewhat tricky to learn proper insights.
@@ D.3 Qualitative Results
Figure 7: Qualitative analysis of object feature space. (a) Class‑level cosine similarity of Ours, (b) is for same as that of VL-SAT [48].